OpenStudy (anonymous):

Complete the following steps to solve the system of linear equations: 3x+4y=18 and x+3y=11. a) Step 1. Make a table of values for each linear equation. Include at least three rows of the table. Be sure to label each table. as in the x input and y output then step 2 tell me the points for this so i can graph

OpenStudy (jiteshmeghwal9):

3x+4y=18 3x=18-4y x=18-4y/3

OpenStudy (anonymous):

what are point for the graph

OpenStudy (anonymous):

points

OpenStudy (anonymous):

help help help help help help help help help help

OpenStudy (anonymous):

can someon please help me im in the middle of a test and need help now

OpenStudy (neo92):

wait bro..

OpenStudy (anonymous):

k

OpenStudy (neo92):

you can simply find two points by substituting y=0 and x=0

OpenStudy (anonymous):

can you do it for me

OpenStudy (neo92):

ok amigo lets consider the equation \[3x+4y=18\] when x=0 \[4y = 18\]\[y=\frac{18}4 = 4.5\] therefore we have the point \[(0,4.5)\]when y =0 \[3x=18\]\[x=\frac{18}3 = 6\]therefore we have the point \[(6,0)\]

OpenStudy (anonymous):

i got the work not just tell me points only

OpenStudy (neo92):

for thre other eqn (0,11/3) (11,0)

OpenStudy (neo92):

is tht wht u wnt u wanna knw tha set of values for different x n y

OpenStudy (anonymous):

now solve the system of linear equations

OpenStudy (anonymous):

hello neo92 i need help now please solve system of linear equations

OpenStudy (neo92):

oh bro sry 4 da l8. did u pot the graphs?

OpenStudy (anonymous):

i now need to solve sytem of equations

OpenStudy (neo92):

solutions are \[x = 2\] and\[y = 3\]

OpenStudy (neo92):

is tht wht u wnt? plot the two graphs and luk for the intersecting point also u can solve these as two simultaneous eqns